# On a B-field Transform of Generalized Complex Structures over Complex Tori

### v1

@@ 2 Gerbes in the sense of Hitchin-Chatterjee
Roughly speaking, for a given smooth manifold $M$, a gerbe on $M$ is determined by an open covering $\{U_{i}\}_{i\in I}$ of $M$ and an element of $H^{2}(M;\mathcal{A}^{*})$, where $\mathcal{A}^{*}$ is the sheaf of nowhere vanishing smooth functions on $M$, so we need to focus on each intersection $U_{ijk}$ of three open sets $U_{i}$, $U_{j}$, $U_{k}$ if we consider gerbes according to the original definition (we use the notations $U_{ij}:=U_{i}\cap U_{j}$, $U_{ijk}:=U_{i}\cap U_{j}\cap U_{k}$, etc. in the above sense). On the other hand, Hitchin and Chatterjee propose a way of defining gerbes as families of line bundles defined on intersections $U_{ij}$ of two open sets. Precisely speaking, this definition is given as follows.
@@
so $\{\Omega_{ij}\}_{i,j\in I}$ defines a representative in $H^{1}(M;\mathcal{A}^{2})$, where $\mathcal{A}^{2}$ is the sheaf of smooth 2-forms on $M$. This implies that the existence of $\beta_{i}\in\Gamma(U_{i};\mathcal{A}^{2})$ such that
@@
We call the family $\beta:=\{\beta_{i}\}_{i\in I}$ of 2-forms $\beta_{i}$ in the above a 1-connection over $\mathcal{G}(I,L,\theta)$ compatible with the 0-connection normal-$\nabla$.
@@ 3 A mirror pair $(X^{n},\check{X}^{n})$
ˇ$\check{\mathcal{I}}:=\left(\begin{array}[]{cccc}I_{n}&O&O&O\\ O&O&O&I_{n}\\ O&O&I_{n}&O\\ O&I_{n}&O&O\end{array}\right)\mathcal{I}\left(\begin{array}[]{cccc}I_{n}&O&O&O\\ O&O&O&I_{n}\\ O&O&I_{n}&O\\ O&I_{n}&O&O\end{array}\right),$ (7)
@@
We can actually rewrite the representation matrix in (12) to
@@
is often called the B-field. Moreover, we define two matrices $\omega_{\rm mat}^{\vee}$ and $B_{\rm mat}^{\vee}$ by
$\omega_{\rm mat}^{\vee}:=\mathrm{Im}(-(T^{-1})^{t})$
@@
Hereafter, for simplicity, we denote
@@
and denote $O_{m_{1}\cdots m_{n}}^{l_{1}\cdots l_{n}}$ by $O_{m}^{l}$ for simplicity:
@@
Namely, when we set
$I:=\Bigl{\{}(l;m)=(l_{1}\cdots l_{n};m_{1}\cdots m_{n})\ |\ l_{1},\cdots,l_{n},m_{1},\cdots,m_{n}=1,2,3\Bigr{\}},$
$\{O_{m}^{l}\}_{(l;m)\in I}$ gives an open covering of $X^{n}$:
$X^{n}=\bigcup_{(l;m)\in I}O_{m}^{l}.$
@@
$L_{mm^{\prime}m^{\prime\prime}}^{ll^{\prime}l^{\prime\prime}}:=\left.L_{mm^{\prime}}^{ll^{\prime}}\right|_{O_{mm^{\prime}m^{\prime\prime}}^{ll^{\prime}l^{\prime\prime}}}\otimes\left.L_{m^{\prime}m^{\prime\prime}}^{l^{\prime}l^{\prime\prime}}\right|_{O_{mm^{\prime}m^{\prime\prime}}^{ll^{\prime}l^{\prime\prime}}}\otimes\left.L_{m^{\prime\prime}m}^{l^{\prime\prime}l}\right|_{O_{mm^{\prime}m^{\prime\prime}}^{ll^{\prime}l^{\prime\prime}}}$
@@
for each $(l;m)\in I$ by using $\tau\in M(n;\mathbb{R})$. Note that this ${}^{\tau}\!B_{m}^{l}$ is defined globally, and we call this globally defined 2-form $\{^{\tau}\!B_{m}^{l}\}_{(l;m)\in I}$ B-field. We consider the B-field transform of the generalized complex structure $\mathcal{I}_{T}$ over $X^{n}$ (see also the local expression (6)) associated to this B-field $\{^{\tau}\!B_{m}^{l}\}_{(l;m)\in I}$:
@@
Now, we define a flat connection over $\mathcal{G}(I,\mathcal{O},\xi)$ by using the matrix $\tau\in M(n;\mathbb{Z})$ which determines the B-field $\{^{\tau}\!B_{m}^{l}\}_{(l;m)\in I}$ as follows. For each $(l;m)$, $(l^{\prime};m^{\prime})\in I$, we take a 1-form ${}^{\tau}\!\omega_{mm^{\prime}}^{ll^{\prime}}$ on the open set $O_{mm^{\prime}}^{ll^{\prime}}$ which is expressed locally as
@@
$\tau_{j}:=\left(\begin{array}[]{ccc}\tau_{j1}\\ \vdots\\ \tau_{jn}\end{array}\right)\in\mathbb{Z}^{n},$
and assume the condition ${}^{\tau}\!\omega_{m^{\prime}m}^{l^{\prime}l}=-^{\tau}\!\omega_{mm^{\prime}}^{ll^{\prime}}$. By using them, we locally define a flat connection ${}^{\tau}\!\nabla_{mm^{\prime}}^{ll^{\prime}}$ on the trivial line bundle $\mathcal{O}_{mm^{\prime}}^{ll^{\prime}}\rightarrow O_{mm^{\prime}}^{ll^{\prime}}$ by
${}^{\tau}\!\nabla_{mm^{\prime}}^{ll^{\prime}}:=d-\mathbf{i}^{\tau}\!\omega_{mm^{\prime}}^{ll^{\prime}}$
@@
over $\mathcal{G}(I,\mathcal{O},\xi)$ compatible with the 0-connection $\nabla_{\tau}$. We denote the trivial gerbe $\mathcal{G}(I,\mathcal{O},\xi)$ with the flat connection $(\nabla_{\tau},B_{\tau})$ by $\mathcal{G}_{\tau}^{\nabla}$:
@@
is the B-field. In particular, associated to the deformation of $X^{n}$ by $\mathcal{G}_{\tau}^{\nabla}$, in the symplectic geometry side, the symplectic form $\omega^{\vee}$ is preserved and the B-field $B^{\vee}$ on $\check{X}^{n}$ is twisted by $\pi d\check{x}^{t}\tau d\check{x}$ (see also the definition (13)). Hereafter, we denote this complexified symplectic torus (the mirror partner of $X_{\mathcal{G}_{\tau}}^{n}$ which is obtained by the definition (7)) by
@@ 5 Deformations of objects and the homological mirror symmetry
In general, the mirror pair $(X^{n},\check{X}^{n})$ can be regarded as the trivial special Lagrangian torus fibrations $X^{n}\rightarrow\mathbb{R}^{n}/\mathbb{Z}^{n}$ and $\check{X}^{n}\rightarrow\mathbb{R}^{n}/\mathbb{Z}^{n}$ on the same base space $\mathbb{R}^{n}/\mathbb{Z}^{n}$ by SYZ construction [34]. Then, we can construct a holomorphic line bundle with an integrable connection on $X^{n}$ from each pair of a Lagrangian section of $\check{X}^{n}\rightarrow\mathbb{R}^{n}/\mathbb{Z}^{n}$ and a unitary local system along it, and this transformation is called SYZ transform (see [28, 2]). Roughly speaking, although the purpose of this section is to discuss deformations of them over the deformed mirror pair $(X_{\mathcal{G}_{\tau}^{\nabla}}^{n},\check{X}_{\mathcal{G}_{\tau}^{\nabla}}^{n})$, when we consider such deformations, it becomes necessary to reconsider how to define objects which should be treated on the deformed complexified symplectic torus under the assumption $\tau\not\in M(n;\mathbb{Z})$ (we deal with this problem in section 6) though all deformations are successfully discussed in the complex geometry side whether the condition $\tau\in M(n;\mathbb{Z})$ is assumed or not. Thus, we assume $\tau\in M(n;\mathbb{Z})$ in subsection 5.3 including the discussions in the symplectic geometry side (we do not assume this condition in subsection 5.2 which is devoted to the discussions in the complex geometry side).
@@
First, we explain the complex geometry side, namely, define a class of holomorphic line bundles $E_{(s,a,q)}\to X^{n}$ with integrable connections $\nabla_{(s,a,q)}$, and hereafter, for simplicity, we sometimes denote a holomorphic line bundle $E_{(s,a,q)}\to X^{n}$ with an integrable connection $\nabla_{(s,a,q)}$ by $E_{(s,a,q)}^{\nabla}$, i.e., $E_{(s,a,q)}^{\nabla}:=(E_{(s,a,q)},\nabla_{(s,a,q)})$. In fact, we first construct it as a smooth complex line bundle on $X^{n}$ with a connection, and discuss the holomorphicity of such a smooth complex line bundle later (Proposition 5.1). Before giving the strict definition of $E_{(s,a,q)}^{\nabla}$, we mention the idea of the construction of $E_{(s,a,q)}$ since the notations of transition functions of it are very complicated. Although we will give the details of the symplectic geometry side again later, in general, the Lagrangian submanifold in $\check{X}^{n}$ corresponding to a holomorphic line bundle with an integrable connection $E_{(s,a,q)}^{\nabla}$ is expressed locally as
@@
$\nabla_{(s,a,q)}=d+\omega_{(s,a,q)},\ \ \ \omega_{(s,a,q)}:=-2\pi\mathbf{i}\Bigl{(}s(x)^{t}+q^{t}T\Bigr{)}dy.$ (17)
@@
We can easily check that the above $\nabla_{(s,a,q)}$ is compatible with the transition functions of $E_{(s,a,q)}$ since the function $s$ satisfies the relations (16). Now, we give the following proposition without its proof since it can be proved similarly as in [23, Proposition 3.1] (see also [21, Proposition 4.3], [22, Proposition 2.1] and [24, Proposition 4.1]), where
@@
Next, we explain the symplectic geometry side, namely, give the objects in the symplectic geometry side corresponding to holomorphic line bundles with integrable connections $E_{(s,a,q)}^{\nabla}$. Let us consider the $n$-dimensional submanifold
@@
$\nabla_{\mathcal{L}_{(s,a,q)}}:=d-2\pi\mathbf{i}q^{t}d\check{x}$ (19)
@@
Hereafter, for simplicity, we set $\mathcal{L}_{(s,a,q)}^{\nabla}:=(\mathcal{L}_{(s,a,q)},\nabla_{\mathcal{L}_{(s,a,q)}})$, and let us denote a pair of an $n$-dimensional submanifold $L_{(s,a)}$ in $\check{X}^{n}$ and the trivial complex line bundle with a flat connection $\mathcal{L}_{(s,a,q)}^{\nabla}$ by $\mathscr{L}_{(s,a,q)}^{\nabla}:=(L_{(s,a)},\mathcal{L}_{(s,a,q)}^{\nabla})$. Then, the following proposition holds. Here, we omit its proof since it can be proved in a similar way which is written in subsection 4.1 in [23] (see also section 4 in [21], section 2 in [22] and [24, Proposition 4.2]).
@@
We explain the interpretation for the above discussions from the viewpoint of SYZ construction. We can regard the complexified symplectic torus $\check{X}^{n}$ as the trivial special Lagrangian torus fibration $\check{X}^{n}\rightarrow\mathbb{R}^{n}/\mathbb{Z}^{n}$, where $\check{x}$ is the local coordinate system of the base space $\mathbb{R}^{n}/\mathbb{Z}^{n}$ and $\check{y}$ is the local coordinate system of the fiber of $\check{X}^{n}\rightarrow\mathbb{R}^{n}/\mathbb{Z}^{n}$. Then, we can regard each affine Lagrangian submanifold $L_{(s,a)}$ in $\check{X}^{n}$ as the graph of the section $s:\mathbb{R}^{n}/\mathbb{Z}^{n}\rightarrow\check{X}^{n}$ of $\check{X}^{n}\rightarrow\mathbb{R}^{n}/\mathbb{Z}^{n}$.

### v2

@@ 2 Gerbes in the sense of Hitchin-Chatterjee
Roughly speaking, for a given smooth manifold $M$, a gerbe on $M$ is determined by an open covering $\{U_{i}\}_{i\in I}$ of $M$ and an element of $H^{2}(M;\mathcal{A}^{*})$, where $\mathcal{A}^{*}$ is the sheaf of nowhere vanishing smooth functions on $M$, so we need to focus on each intersection $U_{ijk}$ of three open sets $U_{i}$, $U_{j}$, $U_{k}$ if we consider gerbes according to the original definition (we use the notations $U_{ij}:=U_{i}\cap U_{j}$, $U_{ijk}:=U_{i}\cap U_{j}\cap U_{k}$, etc. in the above sense). On the other hand, Hitchin and Chatterjee propose a way of defining gerbes as families of line bundles defined on intersections $U_{ij}$ of two open sets. This definition is given as follows.
@@
so $\{\Omega_{ij}\}_{i,j\in I}$ defines a representative in $H^{1}(M;\mathcal{A}^{2})$, where $\mathcal{A}^{2}$ is the sheaf of smooth 2-forms on $M$. This implies the existence of $\beta_{i}\in\Gamma(U_{i};\mathcal{A}^{2})$ such that
@@
We call the family $\beta:=\{\beta_{i}\}_{i\in I}$ of 2-forms $\beta_{i}$ in the above a 1-connection over $\mathcal{G}(I,L,\theta)$ which is compatible with the 0-connection $\nabla$.
@@ 3 A mirror pair $(X^{n},\check{X}^{n})$
ˇ$\check{\mathcal{I}}=\left(\begin{array}[]{cccc}I_{n}&O&O&O\\ O&O&O&I_{n}\\ O&O&I_{n}&O\\ O&I_{n}&O&O\end{array}\right)\mathcal{I}\left(\begin{array}[]{cccc}I_{n}&O&O&O\\ O&O&O&I_{n}\\ O&O&I_{n}&O\\ O&I_{n}&O&O\end{array}\right),$ (7)
@@
We can rewrite the representation matrix in (12) to
@@
is called a B-field. Moreover, we define two matrices $\omega_{\rm mat}^{\vee}$ and $B_{\rm mat}^{\vee}$ by
$\omega_{\rm mat}^{\vee}=\mathrm{Im}(-(T^{-1})^{t})$
@@
Hereafter, we denote this complexified symplectic torus by
@@
and denote $O_{m_{1}\cdots m_{n}}^{l_{1}\cdots l_{n}}$ by $O_{m}^{l}$:
@@
The family $\{O_{m}^{l}\}_{(l;m)\in I}$ of these open sets $O_{m}^{l}$ gives an open covering of $X^{n}$, i.e.,
$X^{n}=\bigcup_{(l;m)\in I}O_{m}^{l},$
where
$I:=\Bigl{\{}(l;m)=(l_{1}\cdots l_{n};m_{1}\cdots m_{n})\ |\ l_{1},\cdots,l_{n},m_{1},\cdots,m_{n}=1,2,3\Bigr{\}}.$
@@
$L_{mm^{\prime}m^{\prime\prime}}^{ll^{\prime}l^{\prime\prime}}=\left.L_{mm^{\prime}}^{ll^{\prime}}\right|_{O_{mm^{\prime}m^{\prime\prime}}^{ll^{\prime}l^{\prime\prime}}}\otimes\left.L_{m^{\prime}m^{\prime\prime}}^{l^{\prime}l^{\prime\prime}}\right|_{O_{mm^{\prime}m^{\prime\prime}}^{ll^{\prime}l^{\prime\prime}}}\otimes\left.L_{m^{\prime\prime}m}^{l^{\prime\prime}l}\right|_{O_{mm^{\prime}m^{\prime\prime}}^{ll^{\prime}l^{\prime\prime}}}$
@@
for each $(l;m)\in I$ by using $\tau\in M(n;\mathbb{R})$. Note that this ${}^{\tau}\!B_{m}^{l}$ is defined globally, and we call this globally defined 2-form $\{^{\tau}\!B_{m}^{l}\}_{(l;m)\in I}$ a B-field. We consider the B-field transform of the generalized complex structure $\mathcal{I}_{T}$ over $X^{n}$ (see also the local expression (6)) associated to this B-field $\{^{\tau}\!B_{m}^{l}\}_{(l;m)\in I}$:
@@
Now, we define a flat connection over $\mathcal{G}(I,\mathcal{O},\xi)$ by using the matrix $\tau\in M(n;\mathbb{R})$ which determines the B-field $\{^{\tau}\!B_{m}^{l}\}_{(l;m)\in I}$ as follows. For each $(l;m)$, $(l^{\prime};m^{\prime})\in I$, we take a 1-form ${}^{\tau}\!\omega_{mm^{\prime}}^{ll^{\prime}}$ on the open set $O_{mm^{\prime}}^{ll^{\prime}}$ which is expressed locally as
@@
$\tau_{j}:=\left(\begin{array}[]{ccc}\tau_{j1}\\ \vdots\\ \tau_{jn}\end{array}\right)\in\mathbb{R}^{n},$
and assume the condition ${}^{\tau}\!\omega_{m^{\prime}m}^{l^{\prime}l}=-^{\tau}\!\omega_{mm^{\prime}}^{ll^{\prime}}$. By using them, we locally define a flat connection ${}^{\tau}\!\nabla_{mm^{\prime}}^{ll^{\prime}}$ on the trivial complex line bundle $\mathcal{O}_{mm^{\prime}}^{ll^{\prime}}\rightarrow O_{mm^{\prime}}^{ll^{\prime}}$ by
${}^{\tau}\!\nabla_{mm^{\prime}}^{ll^{\prime}}=d-\mathbf{i}^{\tau}\!\omega_{mm^{\prime}}^{ll^{\prime}}$
@@
over $\mathcal{G}(I,\mathcal{O},\xi)$ which is compatible with the 0-connection $\nabla_{\tau}$. We denote the trivial gerbe $\mathcal{G}(I,\mathcal{O},\xi)$ with the flat connection $(\nabla_{\tau},B_{\tau})$ by $\mathcal{G}_{\tau}^{\nabla}$:
@@
is a B-field. In particular, associated to the deformation of $X^{n}$ by $\mathcal{G}_{\tau}^{\nabla}$, in the symplectic geometry side, the symplectic form $\omega^{\vee}$ is preserved and the B-field $B^{\vee}$ on $\check{X}^{n}$ is twisted by $\pi d\check{x}^{t}\tau d\check{x}$ (see also the definition (13)). Hereafter, we denote this complexified symplectic torus (the mirror partner of $X_{\mathcal{G}_{\tau}}^{n}$ which is obtained by the definition (7)) by
@@ 5 Deformations of objects and the homological mirror symmetry
In general, the mirror pair $(X^{n},\check{X}^{n})$ can be regarded as the trivial special Lagrangian torus fibrations $X^{n}\rightarrow\mathbb{R}^{n}/\mathbb{Z}^{n}$ and $\check{X}^{n}\rightarrow\mathbb{R}^{n}/\mathbb{Z}^{n}$ on the same base space $\mathbb{R}^{n}/\mathbb{Z}^{n}$ by SYZ construction [35]. Then, we can construct a holomorphic line bundle with an integrable connection on $X^{n}$ from each pair of a Lagrangian section of $\check{X}^{n}\rightarrow\mathbb{R}^{n}/\mathbb{Z}^{n}$ and a unitary local system along it, and this transformation is called SYZ transform (see [29, 2]). The purpose of this section is to discuss deformations of them over the deformed mirror pair $(X_{\mathcal{G}_{\tau}^{\nabla}}^{n},\check{X}_{\mathcal{G}_{\tau}^{\nabla}}^{n})$.
@@
First, we explain the complex geometry side, namely, define a class of holomorphic line bundles $E_{(s,a,q)}\to X^{n}$ with integrable connections $\nabla_{(s,a,q)}$. Hereafter, we sometimes denote a holomorphic line bundle $E_{(s,a,q)}\to X^{n}$ with an integrable connection $\nabla_{(s,a,q)}$ by $E_{(s,a,q)}^{\nabla}$, i.e., $E_{(s,a,q)}^{\nabla}:=(E_{(s,a,q)},\nabla_{(s,a,q)})$. In fact, we first construct it as a smooth complex line bundle on $X^{n}$ with a connection, and discuss the holomorphicity of such a smooth complex line bundle later (Proposition 5.3). Before giving the strict definition of $E_{(s,a,q)}^{\nabla}$, we mention the idea of the construction of $E_{(s,a,q)}$ since the notations of transition functions of it are very complicated. Although we will give the details of the symplectic geometry side again later, in general, the Lagrangian submanifold in $\check{X}^{n}$ corresponding to a holomorphic line bundle with an integrable connection $E_{(s,a,q)}^{\nabla}$ is expressed locally as
@@
$\nabla_{(s,a,q)}=d+\omega_{(s,a,q)},\ \ \ \omega_{(s,a,q)}=-2\pi\mathbf{i}\Bigl{(}s(x)^{t}+q^{t}T\Bigr{)}dy.$ (17)
@@
We can easily check that the above $\nabla_{(s,a,q)}$ is compatible with the transition functions of $E_{(s,a,q)}$ since the function $s$ satisfies the relations (16).
[…]
Definition 5.1 ([6, Definition 1.4.1]).
A connection $D$ is called integrable if $(D^{(0,1)})^{2}=0$.
Therefore, $(E,D)$ can be considered as a holomorphic vector bundle on $X$ if $D$ is integrable.
Remark 5.2.
In this paper, as an analogue of Definition 5.1, for a twisted connection on an $\alpha$-twisted smooth complex vector bundle (this $\alpha$ denotes the 2-cocycle which defines a given gerbe), we call the twisted connection a twisted integrable connection on the $\alpha$-twisted smooth complex vector bundle if the (0,2)-part of its curvature form vanishes.
Now, we give the following proposition without its proof since it can be proved similarly as in [24, Proposition 3.1] (see also [22, Proposition 4.3], [23, Proposition 2.1] and [25, Proposition 4.1]), where
@@
Next, we explain the symplectic geometry side, namely, give the objects in the symplectic geometry side corresponding to holomorphic line bundles with integrable connections $E_{(s,a,q)}^{\nabla}$.
Before starting main discussions, we recall the definition of objects of the Fukaya category over a given complexified symplectic manifold (see [10, Definition 1.1] for example). Let $M_{\omega,B}:=(M,2\pi\omega-2\pi\mathbf{i}B)$ be a complexified symplectic manifold, where $M$ is a smooth even-dimensional real manifold, $2\pi\omega$ is a symplectic form on $M$, and $2\pi B$ is a B-field on $M$. For this $M_{\omega,B}$, an object of the Fukaya category over $M_{\omega,B}$ consists of a Lagrangian submanifold $L$ in $M_{\omega,B}$, i.e., a submanifold $L$ in $M_{\omega,B}$ which satisfies
[…]
and a smooth complex line bundle $\mathcal{L}\to L$ with a unitary connection $\nabla_{\mathcal{L}}$ such that
$\Omega_{\nabla_{\mathcal{L}}}=\Bigl{.}-2\pi\mathbf{i}B\Bigr{|}_{L},$
where $\Omega_{\nabla_{\mathcal{L}}}$ denotes the curvature form of $\nabla_{\mathcal{L}}$.
Let us consider the $n$-dimensional submanifold
@@
$\nabla_{\mathcal{L}_{(s,a,q)}}:=d-2\pi\mathbf{i}q^{t}d\check{x}$
@@
Hereafter, we set $\mathcal{L}_{(s,a,q)}^{\nabla}:=(\mathcal{L}_{(s,a,q)},\nabla_{\mathcal{L}_{(s,a,q)}})$, and let us denote a pair of an $n$-dimensional submanifold $L_{(s,a)}$ in $\check{X}^{n}$ and the trivial complex line bundle with a flat connection $\mathcal{L}_{(s,a,q)}^{\nabla}$ by $\mathscr{L}_{(s,a,q)}^{\nabla}$: $\mathscr{L}_{(s,a,q)}^{\nabla}:=(L_{(s,a)},\mathcal{L}_{(s,a,q)}^{\nabla})$. Then, the following proposition holds. Here, we omit its proof since it can be proved in a similar way which is written in subsection 4.1 in [24] (see also section 4 in [22], section 2 in [23] and [25, Proposition 4.2]).
@@
We give an interpretation for the above discussions from the viewpoint of SYZ construction. We can regard the complexified symplectic torus $\check{X}^{n}$ as the trivial special Lagrangian torus fibration $\check{X}^{n}\rightarrow\mathbb{R}^{n}/\mathbb{Z}^{n}$, where $\check{x}$ is the local coordinate system of the base space $\mathbb{R}^{n}/\mathbb{Z}^{n}$ and $\check{y}$ is the local coordinate system of the fiber of $\check{X}^{n}\rightarrow\mathbb{R}^{n}/\mathbb{Z}^{n}$. Then, we can regard each affine Lagrangian submanifold $L_{(s,a)}$ in $\check{X}^{n}$ as the graph of the section $s:\mathbb{R}^{n}/\mathbb{Z}^{n}\rightarrow\check{X}^{n}$ of $\check{X}^{n}\rightarrow\mathbb{R}^{n}/\mathbb{Z}^{n}$.